The external dimensions of a closed metallic rectangular box are 16cm by 28cm by 40cm.If the metal is 2cm thick. what are the internal measurement of the box? find the volume of the metal used in making the box?
Answer by ikleyn(52861)

.
The external dimensions are 16 cm, 28 cm and 40 cm.


The internal measurements are 16-2*2 = 12 cm,  28-2*2 = 24 cm  and  40-2*2 = 36 cm.  ANSWER.



The volume of the metal = 16*28*40 - 12*24*36 = 7552 cubic centimeters.   ANSWER.